OpenStudy (anonymous):

Given the function f(x) = 2(x - 1)^2 + 1, indicate the shifts that will effect the location of the vertex. What effect will they have? f(x+3) f(x) + 3 f(3x) 3*f(x) Please help.. OMG. .-.

OpenStudy (anonymous):

is the multiple choice related to this question?

OpenStudy (anonymous):

It's part of the question, I guess. I honestly have no idea. It's an essay question, yet, it has those.

OpenStudy (anonymous):

do you recgonize the vertex from below? y = a(x-h)^2 + k

OpenStudy (anonymous):

No, I honestly don't. .-.

OpenStudy (anonymous):

thats ok, but you do see how the vertex form is similar to the equation you give?

OpenStudy (anonymous):

Yes, I do. They're pretty much the same.

OpenStudy (anonymous):

so what would h equal and what would k equal looking at the two forms?

OpenStudy (anonymous):

k would be + 1 and h - 1?

OpenStudy (ranga):

f(x) = 2(x - 1)^2 + 1 compare it to the general vertex form: y = a(x-h)^2 + k where (h,k) is the vertex. You will notice h = 1, k = 1. So the original function has vertex at (1,1) I will do the first one and you follow similar procedure for the others. f(x) = 2(x - 1)^2 + 1 f(x+3) = 2(x + 3 - 1)^2 + 1 = 2(x + 2)^2 + 1 Compare this to y = a(x-h)^2 + k and you will see h = -2 and y = 1. The new vertex is (-2,1). Compare it to the original vertex at (1,1) and you can see the vertex has been shifted to the left by 3 units. You can try the rest.

OpenStudy (anonymous):

k=1 and h=1, these two numbers are the x and y value of the vertext of the parabola. so it means the vertex of the parabola will be 1 to the right and 1 up from the orgin.

OpenStudy (anonymous):

@ranga thanks for the rest of the steps!

OpenStudy (ranga):

no problem. I hope original poster can solve the rest.

OpenStudy (anonymous):

Oh, Jiminy Cricket.. How did you get -2??

OpenStudy (ranga):

2(x + 2)^2 + 1 = 2(x - (-2))^2 + 1 compare it to: y = a(x-h)^2 + k h = -2, k = 1

OpenStudy (anonymous):

OHHH I see! Because h in the formula is negative, whatever number in it's spot is turned to negative as well?

OpenStudy (ranga):

The general form has (x - h) and not (x + h) and so if we see x + 2 we will have to rewrite it as x - (-2) so we can compare it directly to x -h and say h = -2

OpenStudy (anonymous):

Oh, okay, I get that. Now, I have no idea how to set up the others. Can you help me with them, too? I'm absolutely awful at Algebra. Hard to believe I made it to 12th grade with how bad I am.

OpenStudy (ranga):

answering two other questions and I will stop by in between.

OpenStudy (anonymous):

Okay, I'll wait. I'm pretty patient. ^^;;;

OpenStudy (ranga):

f(x) = 2(x - 1)^2 + 1 f(x) + 3 = 2(x - 1)^2 + 1 + 3 = 2(x - 1)^2 + 4 Compare this to y = a(x-h)^2 + k and you will see h = 1 and y = 4 Originally the vertex was (1,1) and now it is (1,4) So the vertex has been shifted up along the y axis by 3. You should try the third part. Just replace x by 3x in the original function and simplify.

OpenStudy (anonymous):

f(3x) = 2(3x - 1)^2 + 1 Did I set that up right?

OpenStudy (ranga):

Yes. Remember we keep comparing the function to y = a(x-h)^2 + k to find the vertex. f(3x) = 2(3x - 1)^2 + 1 Inside the parenthesis we have (3x-1) but in the general function we have (x-h) So we have to get rid of the 3 in front of x. Factor the 3 out. 2(3x - 1)^2 + 1 = 2 { 3(x -1/3) }^2 + 1 = 2 * 3^2 * (x-1/3)^2 + 1 = 18(x-1/3)^2 + 1 Compare this to y = a(x-h)^2 + k and you will see h = 1/3 and y = 1 Originally the vertex was (1,1) and now it is (1/3,1) So the vertex was shrunk to a third when we go from f(x) to f(3x).

OpenStudy (anonymous):

for f(x+3) , the vertex will shift 3 units to the left for f(x)+3 , the vertex will shift 3 units up for f(3x) , the vertex will shift 2/3 units to the left 3*f(x) , the vertex will shift 2 units up

OpenStudy (ranga):

Give the last one a try yourself.

OpenStudy (anonymous):

3*f(x)? I have no idea, but I'll give it a shot. 3*f(x) = 2(3 * x - 1)^2 + 1 Did I set that up right?? It doesn't look right. If you tell me how to set it up, I'll try it out on my own.

OpenStudy (ranga):

3 * f(x) means 3 * { 2(3x - 1)^2 + 1 } = 6(3x-1)^2 + 3 Vertex is (1,3). Compare it to (1,1) and the vertex has shifted up by 2 units. BTW, for part c) say the vertex shifts 2/3 units to the left instead of saying shrunk to a third when we go from f(x) to f(3x).

OpenStudy (anonymous):

OHHH, okay! So the 3 before the f in 3*f(x) is supposed to be multiplied tot he first number outside the brackets? Ah, I think I get it!

OpenStudy (anonymous):

Thank you for your help!!

OpenStudy (ranga):

you are welcome. You can also try the shortcut method that @mangorox used but it would require memorizing a few things. Here the substitutions were fairly simple and we could see how the vertex shifted each time we shifted the original function.